OpenStudy (anonymous):

Model the function rule y = -2x - 3 with a table of values

OpenStudy (zepp):

Toss some numbers in and find some coordinates lying on this function

OpenStudy (zepp):

E.g.: x = 1 y = -2(1) -3 y = -2-3 y = -5 (1, -5) x = 7 y = -2(7)-3 y = -14-3 y = -17 (7, -17)

OpenStudy (anonymous):

should write it like a table

OpenStudy (zepp):